A firm issued a new series of bonds on January 1, 1992. The bonds were sold at par ($1,000), have a 12 percent coupon, and mature in thirty years. Coupon payments are made semi-annually (on June 30 and December 31).

If, on July 1, 2012, an investor expects the bonds to sell for $896.64. What is the expected yield to maturity on the bonds at that date?

Please help using a financial calculator if possible!
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE1 De Bano VD (dx 0523 ! ENG 220120204 Į K O P Q BA PLUS II calculator strokes FOR YTM first of all set 2 C/Y = P/Y= then N =